Please would someone help me with this math problem?
C is the midpoint of line ED, line CD = 6 + x and line CE = 2x + 1 , what is the length of line ED?
thank you

Answer by fcabanski(1391)

If C is the midpoint, then the distance from C to one end (E) is the same as (equal to) the distance from C to the other end (D).


CE = CD ---> 2x+1 = 6+x


x=5


CE = 2(5)+1 = 11 and CD=6+5=11.


ED = CE + CD = 11+11 = 22